OpenStudy (anonymous):

f(x)=x^2-6x-3 For x I got 3, but I am having trouble finding the y coordinate. Can someone please help?

OpenStudy (johnweldon1993):

for x you got 3? ...well okay remember f(x) is the same as y so y = x² - 6x - 3 just plug your 3 in for the x's y = (3)² - 6(3) - 3 y = 9 - 18 - 3 y = -12

OpenStudy (anonymous):

Ok I had that as far as the equation goes and plugging in 3, clearly my math is off when adding it all together! Thanks so much!

OpenStudy (johnweldon1993):

no problem :)